OpenStudy (anonymous):

Please help? Given the function f(x) = 3(x − 3)2 + 2, indicate the shifts that will affect the location of the vertex, and explain what effect they will have. Use complete sentences. f(4x) 4•f(x)

OpenStudy (anonymous):

the function of f(4x) is 4x times 4 times the 4x. then divide to 2 will still be 4!

OpenStudy (anonymous):

what? i am so confused. >.<

OpenStudy (yanasidlinskiy):

I gotcha:D

OpenStudy (anonymous):

thank you so muchhh

OpenStudy (yanasidlinskiy):

f(x + 4) shifts the function to the left by 4 units, which results in the vertex shifting from (3,2) to (-1,2). f(x) + 4 shifts the function vertically by 4 units, which results in the vertex shifting to (3,6). The transformed function is g(x) = f(4x) = 3*(4x - 3)^2 + 2 = 48*(x - (3/4))^2 + 2, so the vertex shifts to ((3/4), 2). The transformed function is g(x) = 4*f(x) = 12*(x - 3)^2 + 8, so the vertex shifts to (3,8).